OpenStudy (anonymous):

dividing radicals: what's 12/2-√15 is it 2√15/5 or am i wrong? I don't understand wolfram alpha

OpenStudy (anonymous):

it is telling you if you rationalize the denominator you will get \[-\frac{24+12\sqrt{15}}{11}\][

OpenStudy (anonymous):

easier just to do it. write \[\frac{12}{2-\sqrt{15}}\times \frac{2+\sqrt{15}}{2+\sqrt{15}}=\frac{12(2+\sqrt{15})}{-11}\]

OpenStudy (anonymous):

oh i see where i went wrong thanks!